If the length of the sides of the cube is 2.1 feet, which is closest to the surface area of the box?

Answer by colliefan(242)

The area of one side is:
A = l * w
For a square, A = s^2 since the sides are equal.
A=2.1 ft * 2.1 ft
A= 4.41 ft^2
Since there are 6 sides on the cube, the surface area is 6 times the area of one side.
SA = 6 * 4.41 ft^2 = 26.46 ft^2